OpenStudy (geekfromthefutur):

Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points) Part B: How many liters of apple juice are present in the jar? Show your work. (5 points)

OpenStudy (geekfromthefutur):

@SiennaSurfer help me plz

OpenStudy (anonymous):

is this flvs?

OpenStudy (geekfromthefutur):

yea :/

OpenStudy (anonymous):

What grade? and lesson?

OpenStudy (geekfromthefutur):

9th and its 2.10 Module Two Test

OpenStudy (anonymous):

let me look... i did that already

OpenStudy (geekfromthefutur):

ok

OpenStudy (anonymous):

danget... im in the middle of a test myself

OpenStudy (geekfromthefutur):

ah wat question ?

OpenStudy (anonymous):

im on the segment 1 thing

OpenStudy (geekfromthefutur):

oh ok

OpenStudy (anonymous):

ill check once im done... ill work fast

OpenStudy (geekfromthefutur):

ok thax :3

OpenStudy (anonymous):

im on question 21 out of 58... i wont be done for a while

OpenStudy (geekfromthefutur):

oh ok wait are u in algebra or pre algebra

OpenStudy (anonymous):

algebra

OpenStudy (geekfromthefutur):

ohok same here

OpenStudy (geekfromthefutur):

here is could see if someone can help us if u want ?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

question 24

OpenStudy (geekfromthefutur):

@Hero can u help us plz?

hero (hero):

\[\frac{75\%}{25\%} = \frac{x}{1}\]

OpenStudy (anonymous):

25

hero (hero):

\[\frac{75\%}{25\%} = \frac{x \space \text{liters}}{1 \space\text{liters}}\]

OpenStudy (geekfromthefutur):

oh ok is that the answer for part A?

hero (hero):

x = liters of apple juice

OpenStudy (anonymous):

The table below shows two equations. Equation 1 |4x – 3| – 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations? Equation1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = -4, 1. The solutions to equation 1 are x = 3, -1.5 and equation 2 has no solution. The solutions to equation 1 are x = 3, -1.5 and equation 2 has solutions x = -4, 1.

OpenStudy (geekfromthefutur):

i havethe same Q i think hold on really quick

hero (hero):

@siennasurfer, I saw your question you posted in your own area. I was about to help. But don't get into the habit of posting your question in someone else's question. It is generally considered rude to do so.

OpenStudy (anonymous):

we r helping each other. i dont norm do this... but we r working as a team to get these done.

hero (hero):

Okay, it's fine then I suppose.

OpenStudy (geekfromthefutur):

how many questions do u have left ?

OpenStudy (anonymous):

im on 33

OpenStudy (geekfromthefutur):

oh ok is there any questions u need help on ?

OpenStudy (anonymous):

ya... i posted it above

OpenStudy (anonymous):

# 36

OpenStudy (anonymous):

#40

OpenStudy (anonymous):

#45

OpenStudy (geekfromthefutur):

oh ok :)

OpenStudy (anonymous):

# 50

OpenStudy (anonymous):

55

OpenStudy (geekfromthefutur):

oh ok is that all your questions?

OpenStudy (anonymous):

3 more

OpenStudy (anonymous):

2 left

OpenStudy (anonymous):

last 1

OpenStudy (anonymous):

done!

OpenStudy (geekfromthefutur):

what grade did u get on it?

OpenStudy (anonymous):

73

OpenStudy (anonymous):

it was just the practice test

OpenStudy (anonymous):

only worth a point

OpenStudy (geekfromthefutur):

what module are u in?

OpenStudy (anonymous):

I have the segment exams... then im done with the semester

OpenStudy (anonymous):

what do u need help with?

OpenStudy (anonymous):

what module?

OpenStudy (geekfromthefutur):

i was needign help with my module 2 test part 2 :3

OpenStudy (anonymous):

i failed that... but i got #2 completely right

OpenStudy (anonymous):

i dont have much time... goin to bed real soon

OpenStudy (geekfromthefutur):

oh ok u dont mind if i could see the ones u got right

OpenStudy (anonymous):

im not gonna just give u the answers...

OpenStudy (anonymous):

ill help u through # 2 though

OpenStudy (geekfromthefutur):

oh ok

OpenStudy (anonymous):

Jake is comparing the prices of two mattress cleaning companies. Company A charges $30 per mattress and an additional $13 as service charges. Company B charges $28 per mattress and an additional $15 as service charges.

OpenStudy (anonymous):

Part A: Write equations to represent Company A and Company B's total mattress cleaning charges for a certain number of mattresses. Define the variable used in the equations. (4 points)

OpenStudy (anonymous):

thats it... right?

OpenStudy (anonymous):

bthw u should close this conversation

OpenStudy (geekfromthefutur):

i dont have that question :/

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

wait... what course r u takin

OpenStudy (geekfromthefutur):

im taking algebra Algebra I V17

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

be right back

OpenStudy (geekfromthefutur):

ok

OpenStudy (anonymous):

back

OpenStudy (anonymous):

@geekfromthefutur ? u on?

OpenStudy (geekfromthefutur):

yea sorry i was in the kitchen

OpenStudy (anonymous):

me too

OpenStudy (geekfromthefutur):

lol

OpenStudy (anonymous):

:)

OpenStudy (geekfromthefutur):

:) wats the other question u got right

OpenStudy (anonymous):

whats ur first q?

OpenStudy (geekfromthefutur):

Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points) Part B: How many liters of apple juice are present in the jar? Show your work. (5 points)

OpenStudy (anonymous):

this is what i put... part a: .75(1)=a a=apple juice

OpenStudy (geekfromthefutur):

oh ok did the teacher give u a good grade ?

OpenStudy (anonymous):

in my answer, i didnt put what a represented... so in that part, i got 3/5

OpenStudy (geekfromthefutur):

oh ok

OpenStudy (anonymous):

part b: .75(1)=a .75=a .75 liters of apple juice in bottle.

OpenStudy (anonymous):

make sure u show ur work

OpenStudy (geekfromthefutur):

ohok thsnk you :3

OpenStudy (anonymous):

whats #2?

OpenStudy (anonymous):

thx

OpenStudy (geekfromthefutur):

yw (:

OpenStudy (geekfromthefutur):

Maria is comparing the prices of two window cleaning companies. Company A charges $6 per window and an additional $12 as service charges. Company B charges $5 per window and an additional $15 as service charges. Part A: Write equations to represent Company A and Company B's total charges for cleaning a certain number of windows. Define the variable used in the equations. (4 points) Part B: Which company would charge less for cleaning 8 windows? Justify your answer. (3 points) Part C: How much money is saved by using the services of Company B instead of Company A to clean 6 windows? (3 points)

OpenStudy (anonymous):

part a: compony a: 6w+12 compony b: 5w+15

OpenStudy (anonymous):

sorry, gtg. my mom is home. i gotta go to bed. night

OpenStudy (geekfromthefutur):

night thx for the help (: